OpenStudy (anonymous):

Solve the equation. x – = 3 A. x = –3 B. x = –2 C. x = 2 D. x = 3

OpenStudy (anonymous):

x – = 3

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

hello

OpenStudy (qwertty123):

hi

OpenStudy (anonymous):

hi

OpenStudy (qwertty123):

Sorta lost on your question? You want to solve...x-=3. Is that the equation to solve?

OpenStudy (anonymous):

no its x-2/7=3

OpenStudy (qwertty123):

oh okay!

OpenStudy (anonymous):

yea sorry

OpenStudy (qwertty123):

No problemo! :) So you are going to want to simplify both of the equations. x+-2/7=3! Then Step to you would do what?

OpenStudy (anonymous):

not understanding

OpenStudy (qwertty123):

oh okay um you have to simplify it! Then Add 2/7 to both sides!

OpenStudy (anonymous):

?????

OpenStudy (qwertty123):

what is the equation?

OpenStudy (anonymous):

x-2/7=3

OpenStudy (qwertty123):

okay so to simplify it you are doing to need to get x by itself

OpenStudy (anonymous):

clueless

OpenStudy (anonymous):

im sorry

OpenStudy (qwertty123):

no I am just trying to see how I can explain this!!! :)

OpenStudy (qwertty123):